OpenStudy (anonymous):

Make "l" the subject of T = 1/5 sqrt of "l" (Last part after the T is one fifth to the square root of the letter L)

OpenStudy (anonymous):

\[T=\frac{1}{\sqrt{L}}\]

OpenStudy (anonymous):

take the reciprocal of both sides \[\frac{1}{T}=\sqrt{L}\]

OpenStudy (anonymous):

No Sat, one fifth to the square root of l

OpenStudy (anonymous):

then square both sides \[(\frac{1}{T})^2=l\]

OpenStudy (anonymous):

OHHHHH ! Was that an example ? lol

OpenStudy (anonymous):

ok fine l not L

OpenStudy (anonymous):

hold the phone. is it \[T=\frac{1}{\sqrt[5]{l}}\]

OpenStudy (anonymous):

fifth root?

OpenStudy (anonymous):

in that case take the reciprocal of both sides and raise to the fifth power.

OpenStudy (anonymous):

\[\frac{1}{T}=\sqrt[5]{l}\]

OpenStudy (anonymous):

\[(\frac{1}{T})^5=l\] \[\frac{1}{T^5}=l\]

OpenStudy (anonymous):

Sat, ONE FIFTH like the fraction 1 OVER 5.... its \[T = 1/5 \sqrt{l}\]

OpenStudy (anonymous):

ok we try again

OpenStudy (anonymous):

\[T=\frac{1}{5\sqrt{l}}\]

OpenStudy (anonymous):

take the reciprocal \[\frac{1}{T}=5\sqrt{l}\]

OpenStudy (anonymous):

divide both sides by 5 \[\frac{1}{5T}=\sqrt{l}\]

OpenStudy (anonymous):

square both sides \[(\frac{1}{5t})^2=l\] \[\frac{1}{25T^2}=l\]

OpenStudy (anonymous):

Thanks !! Love the explanation.